Find the amount of money that would be created in the banking system because of the money multiplier if the required reserve ratio is 10%, and a bank that had been holding $250 as excess reserves decides to loan all this money out
Answer #1

Answer-

Money multiplier (m) = 1/reserve ratio

=1/10%= 10

Money supply= excess reserve × money multiplier

=$250 × 10 = $2500

Money created in banking system will be $2500.